# On the Identifying Power of Generalized Monotonicity

### v3

@@ Appendix A Proofs of Main Results
The desired result follows immediately from Theorems 3.2 and 3.3.
@@
In Section A.3.2, we use the notation and lemma introduced in Section A.3.1 to prove the theorem. Let $\mathbf{Q}$ satisfy the assumptions of the theorem. We first show that $\Theta_{0}(P,\mathbf{Q})$ is a subset of the bounds in (5). We then show that the bounds in (5) are a subset of $\Theta_{0}(P,\mathbf{Q})$ using the following proof strategy. By assumption, $\Theta_{0}(P,\mathbf{Q})$ is non-empty, so that there exists a distribution $Q\in\mathbf{Q}$ that rationalizes $P$. For each value $\theta_{0}$ in the bounds of (5), we construct an alternative distribution $Q^{\ast}\in\mathbf{Q}$ such that $\theta(Q^{\ast})=\theta_{0}$ with $Q$ and $Q^{\ast}$ differing only in the probability of outcomes corresponding to treatments that a given response type would not take even when maximally encouraged to do so. That the constructed $Q^{\ast}$ lies in $\mathbf{Q}$ follows from the assumption that $\mathbf{Q}|_{Y}$ satisfies Assumption 3.1 and that $Q$ and $Q^{\ast}$ have the same distribution of potential treatment choices with $Q\in\mathbf{Q}$. That the constructed $Q^{\ast}$ rationalizes $P$ follows from $Q\in\mathbf{Q}$ and the previously described lemma. That we are able to construct such a $Q^{\ast}$ with $\theta(Q^{\ast})=\theta_{0}$ for every $\theta_{0}$ in the bounds of (5) establishes that the bounds (5) are a subset of $\Theta_{0}(P,\mathbf{Q})$, completing the proof.
@@ A.3.1 Auxillary Results
Following Heckman and Pinto (2018), we define a treatment response type as a vector $r^{t}\in\mathcal{D}^{|\mathcal{Z}|}$,
@@
Treatment response types are also called principal strata (Frangakis and Rubin, 2002). We analogously define an outcome response type as a vector $r^{o}\in\mathcal{Y}^{|\mathcal{D}|}$,
@@
$\mathcal{N}(r^{t})$ =$=\{d\in\mathcal{D}~{}:~{}r^{t}_{z}\neq d~{}\mbox{for all}~{}z\in\mathcal{Z}\}~{},$
$\mathcal{N}(r^{t})^{c}$ =$=\{d\in\mathcal{D}~{}:~{}r^{t}_{z}=d~{}\mbox{for some}~{}z\in\mathcal{Z}\}~{},$
@@
$\mathcal{N}(r^{t})$ =$=\{d\in\mathcal{D}~{}:~{}r^{t}_{z^{\ast}(d)}\neq d\}~{},$
$\mathcal{N}(r^{t})^{c}$ =$=\{d\in\mathcal{D}~{}:~{}r^{t}_{z^{\ast}(d)}=d\}~{},$
@@
For notational convenience, we further define the probability mass $q(r^{o}_{c}(r^{t}),r^{t})$ as $q(r^{o}_{n}(r^{t}),r^{o}_{c}(r^{t}),r^{t})$ summed over $r^{o}_{n}(r^{t})$:
@@ A.3.2 Proof of the Theorem
    satisfies $\sum_{r^{o}}q^{\ast}(r^{o},r^{t})=\sum_{r^{o}}q(r^{o},r^{t})$ and hence $Q^{\ast}\in\mathbf{Q}$ because the distribution of $(D_{z}:z\in\mathcal{Z})$ is unchanged and by assumption that $\mathbf{Q}|_{Y}$ satisfies Assumption 3.1,
@@
Because we assume $\mathbf{Q}|_{Y}$ satisfies Assumption 3.1 and $Q\in\mathbf{Q}$, this implies that $Q^{\ast}\in\mathbf{Q}$.
@@
Suppose $Q\in\mathbf{Q}_{0}(P,\mathbf{Q}_{E,M}^{\ast})\neq\emptyset$. Consider the upper endpoints of (5) and (8). For any $d\in\mathcal{D}$, $z\in\mathcal{Z}$,
@@
$=\mathbb{E}_{Q}[Y_{d}\mathbbm{1}\{D_{z^{\ast}(d)}=d\}]+y^{U}\mathbb{E}_{Q}[1-\mathbbm{1}\{D_{z^{\ast}(d)}=d\}]-\mathbb{E}_{Q}[Y_{d}\mathbbm{1}\{D_{z}=d\}]-y^{U}\mathbb{E}_{Q}[1-\mathbbm{1}\{D_{z}=d\}]$
@@
[…]
By Lemma 3.2, when $\mathbf{Q}_{0}(P,\mathbf{Q}_{E,M}^{\ast})\neq\emptyset$, we have
[…]
@@ Appendix B Details of Examples
We show that, except in the special case where the treatment and the instrument are both binary, the strict one-to-one targeting assumption of Lee and Salanié (2023) with one or more targeted treatments implies that (11) does not hold for some treatments. To see this, suppose that the strict one-to-one targeting assumption of Lee and Salanié (2023) holds with $|\mathcal{D}^{\dagger}|\geq 1$. For each $d\in\mathcal{D}^{\dagger}$, their assumptions include that there exists some $z^{\dagger}(d)$ and some $\overline{U}(d)$, $\underline{U}(d)$ with $\overline{U}(d)>\underline{U}(d)$ such that
@@
$g(z,d^{\prime})>g(z^{\ast}(d),d^{\prime})~{}~{}\text{ for all }d^{\prime}\neq d\text{ and }z\neq z^{\ast}(d)~{}.$ (20)
@@
We have shown (11) does not hold for some treatments when either $Z$ or $D$ takes at least three values. Now suppose $|\mathcal{D}|=|\mathcal{Z}|=2$. Let $D=0$ denote the nontargeted treatment and $D=1$ the targeted treatment, and let $z^{\dagger}(1)=1$. Consider $z^{\ast}(0)=0$ and $z^{\ast}(1)=1$. Then evaluating (11) at either $d=0$ or $d=1$ results in $\overline{U}(1)>\underline{U}(1),$ and thus (11) holds when $|\mathcal{D}|=|\mathcal{Z}|=2$. We conclude that the strict one-to-one targeting of Lee and Salanié (2023) implies that (11) does not hold for some $d\in\mathcal{D}$ except in the special case where $|\mathcal{D}|=|\mathcal{Z}|=2$.
We now show that the strict one-to-one targeting of Lee and Salanié (2023) implies that Assumption 2.2 holds when $|\mathcal{Z}|>|\mathcal{D}^{\dagger}|$. Let $\mathcal{Z}^{\dagger}\subseteq\mathcal{Z}$ denote the set of instruments that target some treatment,
@@
Their strict one-to-one targeting assumption combined with $|\mathcal{Z}|>|\mathcal{D}^{\dagger}|$ implies that there are values of the instrument that do not target any treatment; in other words, $\mathcal{Z}^{\dagger}\subsetneqq\mathcal{Z}$. Following Lee and Salanié (2023), we label the treatment that is known not to be targeted as treatment 00, so that $g(z,0)=\underline{U}(0)$ for all $z\in\mathcal{Z}$, and impose their normalization that $\underline{U}(0)=0$. Consider (3) for $d=0$. Note that
@@
Thus, the event $\{D_{z^{\ast}(0)}=d^{\ast},~{}D_{z^{\prime}}=0\}$ either requires (21) and (22) to jointly hold, which is a contradiction since $\overline{U}(d^{\ast})>\underline{U}(d^{\ast})$ , or requires (21) and (23) to jointly hold, which is a zero probability event given our assumption that the distribution of $(U_{d}:d\in\mathcal{D})$ is absolutely continuous w.r.t. Lebesgue measure. Thus $Q\{D_{z^{\ast}(0)}\neq 0,~{}D_{z^{\prime}}=0~{}\mbox{for some}~{}z^{\prime}\neq z^{\ast}(0)\}$ is a probability of a finite union of zero probability events, and thus, by Boole’s inequality, equals zero so that (3) holds for $d=0$. A parallel argument shows that (3) holds for any non-targeted treatment, and related argument shows that (3) holds for any targeted treatment. Thus, under the strict one-to-one targeting of Lee and Salanié (2023), when there are more values of the instrument than targeted treatments, Assumption 2.2 holds even though (11) is violated for some $d\in\mathcal{D}$.
@@
In contrast, the identified set that follows from imposing Assumption 2.1 alone, $\Theta_{0}(P,\mathbf{Q}_{E}^{\ast})$, is shown by Balke and Pearl (1997) to be
@@
Suppose $\mathcal{Y}=\{0,1\}$, $\mathcal{D}=\{0,1,2\}$, and $\mathcal{Z}=\{0,1,2\}$. Then, the linear program approach in Balke and Pearl (1993, 1997) leads to the following identified set for $\mathbb{E}_{Q}[Y_{1}]=Q\{Y_{1}=1\}$ relative to $\mathbf{Q}$ being defined as in Example 5.2:
@@
Consider the ARUM defined by (10) with strict one-to-one targeting and $|\mathcal{D}|=3$, $|\mathcal{Z}|=|\mathcal{D}^{\dagger}|=2$. In this case, there are two targeted treatments and one non-targeted treatment. Following Lee and Salanié (2023), label that non-targeted treatment as treatment 00 and impose the normalization that $g(z,0)=0$ for all $z\in\mathcal{Z}$. Label $Z=0$ as the instrument value that targets treatment $1$ and label $Z=1$ as the instrument value that targets treatment $2$, so that (19) holds for $d=1,2$ for some $\overline{U}(d),\underline{U}(d)$ with $\overline{U}(d)>\underline{U}(d)$ and with $z^{\dagger}(1)=0$, $z^{\dagger}(2)=1$. Let $\mathbf{Q}$ denote the set of all distributions for which $(D_{z}:z\in\mathcal{Z})$ is determined by (10) with these restrictions and additionally imposing that the support of $(U_{0},U_{1},U_{2})$ is $\Re^{3}$. Recall $(U_{0},U_{1},U_{2})\perp\!\!\!\perp Z$ by assumption.
@@
and a similar characterization holds for $D_{z}=1$. See Figure 1, which is taken from Figure 1 in Lee and Salanié (2023).
@@
We now show that there does not exist a value of $z^{\ast}(0)$ such that (3) holds for the non-targeted treatment, treatment 00. Suppose $z^{\ast}(0)=0$. Then
@@
where the last line is using that the support of the distribution of $(U_{10},U_{20})=\Re^{2}$ by assumption and that strict targeting of treatment $1$ requires $-\underline{U}(1)>-\overline{U}(1)$. Thus (3) cannot hold for $d=0$ with $z^{\ast}(0)=0$. A parallel argument shows that (3) cannot hold for $d=0$ with $z^{\ast}(0)=1$.
@@
We now derive the identified set for the average potential outcome of the non-targeted treatment. First, note that $-g(0,1)=-\overline{U}(1)<-\underline{U}(1)=-g(1,1)$ and $-g(0,2)=-\underline{U}(2)>-\overline{U}(2)=-g(1,2)$. Therefore, it can be verified from Figure 2 that for all $Q\in\mathbf{Q}$,
@@
Let $\mathbf{Q}^{\prime}$ denote the set of all distributions that satisfies (29). Note that all $Q\in\mathbf{Q}$ satisfies (29), so $\mathbf{Q}\subseteq\mathbf{Q}^{\prime}$. On the other hand, by assigning appropriate probabilities to each set in the partition in Figure 2, we immediately see that each $Q\in\mathbf{Q}^{\prime}$ can be rationalized by a $Q\in\mathbf{Q}$. Therefore, $\mathbf{Q}=\mathbf{Q}^{\prime}$. Using linear programming as in Balke and Pearl (1993, 1997), we obtain the following identified set for $\mathbb{E}_{Q}[Y_{0}]=Q\{Y_{0}=1\}$ relative to $\mathbf{Q}$:
@@
Finally, we show that there exists a $P$ for which $\mathbf{Q}_{0}(P,\mathbf{Q})\neq\emptyset$ and (30) is strictly smaller than (5), so that $\Theta_{0}(P,\mathbf{Q})$ is not given by (5). We do so by providing a numerical example. Consider the $P$ specified in Table 8 and the $Q_{\text{arum},\text{min}}$ and $Q_{\text{arum},\text{max}}$ specified in Tables 9 and 10 respectively, where we write $q(y_{0}y_{1}y_{2},d_{0}d_{1})=Q\{Y_{d}=y_{d},D_{z}=d_{z},~{}(d,z)\in\mathcal{D}\times\mathcal{Z}\}$ and omit any $q(\cdot)=0$. One can check that both $Q_{\text{arum},\text{min}}$ and $Q_{\text{arum},\text{max}}$ rationalize $P$ and satisfy Assumption 2.1. One can further check that both $Q_{\text{arum},\text{min}}$ and $Q_{\text{arum},\text{max}}$ satisfy the restriction in (29), so that $\mathbf{Q}_{0}(P,\mathbf{Q})\neq\emptyset$. Evaluating (30) at $P$ gives the identified set for $\mathbb{E}_{Q}[Y_{0}]$ relative to $\mathbf{Q}$ as $[0.2518,0.8167]$. One can further check that the two endpoints are attained by $\mathbb{E}_{Q_{\text{arum},\text{min}}}[Y_{0}]=0.2518$ and $\mathbb{E}_{Q_{\text{arum},\text{max}}}[Y_{0}]=0.8167$. On the other hand, if one evaluates (5) by setting $z^{\ast}(0)=0,1$ at the same $P$, the resulting bounds for $\mathbb{E}_{Q}[Y_{0}]$ equal $[p_{10|0},1-p_{00|0}]=[0.2518,0.8937]$ and $[p_{10|1},1-p_{00|1}]=[0.2372,0.8167]$ respectively. In both cases, (30) is strictly contained in (5). Hence, $\Theta_{0}(P,\mathbf{Q})$ is not given by (5).
@@
Kline and Walters (2016) considers an RCT with a “close substitute” to study the effects of preschooling on educational outcomes. In their setting, $D\in\mathcal{D}=\{0,1,2\}$, where $D=0$ denotes home care (no preschool), $D=2$ denotes a preschool program called Head Start, and $D=1$ denotes preschools other than Head Start, namely the close substitute. Let $Z\in\mathcal{Z}=\{0,1\}$ denote an indicator variable for an offer to attend Head Start. Assumption 2.1 holds because $Z$ is randomly assigned. Kline and Walters (2016) impose the restriction that
@@
Note in this example Assumption 2.2 still holds although $|\mathcal{Z}|<|\mathcal{D}|$. See Bai et al. (2025) for results on the sharp testable implications of the assumptions for this example and Example C.3.
@@
Kirkeboen et al. (2016) study the effects of fields of study on earnings. In their setting, $\mathcal{D}=\{0,1,2\}$ represent three fields of study, ordered by their (soft) admission cutoffs from the lowest to the highest. The instrument is $Z\in\{0,1,2\}$, with $Z=1$ when the student crosses the (soft) admission cutoff for field 1, $Z=2$ when the student crosses the (soft) admission cutoff for field 2, and $Z=0$ otherwise. The authors assume that $Z$ is exogenous in the sense that $Q$ satisfies Assumption 2.1 and impose the following monotonicity conditions:

### v4

@@ Appendix A Proofs of Main Results
The desired result follows immediately from Theorems 3.2 and $3.$
@@
In Section A.3.2, we use the notation and lemma introduced in Section A.3.1 to prove the theorem. Let $\mathbf{Q}$ satisfy the assumptions of the theorem. We first show that $\Theta_{0}(P,\mathbf{Q})$ is a subset of the bounds in (6). We then show that the bounds in (6) are a subset of $\Theta_{0}(P,\mathbf{Q})$ using the following proof strategy. By assumption, $\Theta_{0}(P,\mathbf{Q})$ is non-empty, so that there exists a distribution $Q\in\mathbf{Q}$ that rationalizes $P$. For each value $\theta_{0}$ in the bounds of (6), we construct an alternative distribution $Q^{\ast}\in\mathbf{Q}$ such that $\theta(Q^{\ast})=\theta_{0}$ with $Q$ and $Q^{\ast}$ differing only in the probability of outcomes corresponding to treatments that a given response type would not take even when maximally encouraged to do so. That the constructed $Q^{\ast}$ lies in $\mathbf{Q}$ follows from the assumption that $\mathbf{Q}$ satisfies Assumption 3.1 and that $Q$ and $Q^{\ast}$ have the same distribution of potential treatment choices with $Q\in\mathbf{Q}$. That the constructed $Q^{\ast}$ rationalizes $P$ follows from $Q\in\mathbf{Q}$ and the previously described lemma. That we are able to construct such a $Q^{\ast}$ with $\theta(Q^{\ast})=\theta_{0}$ for every $\theta_{0}$ in the bounds of (6) establishes that the bounds (6) are a subset of $\Theta_{0}(P,\mathbf{Q})$, completing the proof.
@@ A.3.1 Auxillary Results
Following heckman2018unordered, we define a treatment response type as a vector $r^{t}\in\mathcal{D}^{|\mathcal{Z}|}$,
@@
Treatment response types are also called principal strata (frangakis2002principal). We analogously define an outcome response type as a vector $r^{o}\in\mathcal{Y}^{|\mathcal{D}|}$,
@@
$\mathcal{N}(r^{t})$ $=\{d\in\mathcal{D}~:~r^{t}_{z}\neq d~\mbox{for all}~z\in\mathcal{Z}\}~,$
$\mathcal{N}(r^{t})^{c}$ $=\{d\in\mathcal{D}~:~r^{t}_{z}=d~\mbox{for some}~z\in\mathcal{Z}\}~,$
@@
$\mathcal{N}(r^{t})$ $=\{d\in\mathcal{D}~:~r^{t}_{z^{\ast}(d)}\neq d\}~,$
$\mathcal{N}(r^{t})^{c}$ $=\{d\in\mathcal{D}~:~r^{t}_{z^{\ast}(d)}=d\}~,$
@@
For notational convenience, we further define the probability mass $q(r^{o}_{c}(r^{t}),r^{t})$ as $q(r^{o}_{n}(r^{t}),\allowbreak r^{o}_{c}(r^{t}),r^{t})$ summed over $r^{o}_{n}(r^{t})$:
@@ A.3.2 Proof of the Theorem
    satisfies $\sum_{r^{o}}q^{\ast}(r^{o},r^{t})=\sum_{r^{o}}q(r^{o},r^{t})$ and hence $Q^{\ast}\in\mathbf{Q}$ because the distribution of $(D_{z}:z\in\mathcal{Z})$ is unchanged and by assumption that $\mathbf{Q}$ satisfies Assumption 3.1,
@@
Because we assume $\mathbf{Q}$ satisfies Assumption 3.1 and $Q\in\mathbf{Q}$, this implies that $Q^{\ast}\in\mathbf{Q}$.
@@
Suppose $Q\in\mathbf{Q}_{0}(P,\mathbf{Q})\neq\emptyset$. Consider the upper endpoints of (6) and (9). For any $d\in\mathcal{D}$, $z\in\mathcal{Z}$,
@@
$=\mathbb{E}_{Q}[Y_{d}\mathbbm{1}\{D_{z^{\ast}(d)}=d\}]+y^{U}\mathbb{E}_{Q}[1-\mathbbm{1}\{D_{z^{\ast}(d)}=d\}]$
$\hskip 30.00005pt-\mathbb{E}_{Q}[Y_{d}\mathbbm{1}\{D_{z}=d\}]-y^{U}\mathbb{E}_{Q}[1-\mathbbm{1}\{D_{z}=d\}]$
@@
Because by assumption $\mathbf{Q}\subseteq\mathbf{Q}^{\prime}\subseteq\mathbf{Q}_{\it MI}^{\ast}$, we have
$\Theta_{0}(P,\mathbf{Q})\subseteq\Theta_{0}(P,\mathbf{Q}^{\prime})\subseteq\Theta_{0}(P,\mathbf{Q}_{\it MI}^{\ast})~.$
By Lemma 3.2, if $\mathbf{Q}\subseteq\mathbf{Q}_{E,M}^{\ast}$, $\mathbf{Q}$ satisfies Assumption 3.1, and $\mathbf{Q}_{0}(P,\mathbf{Q})\neq\emptyset$, then we have
$\Theta_{0}(P,\mathbf{Q})=\Theta_{0}(P,\mathbf{Q}_{\it MI}^{\ast})~.$
@@
A.8 Proof of Corollary 3.2
The result follows by taking $\mathbf{Q}=\mathbf{Q}_{E,M}^{\ast}$ and $\mathbf{Q}^{\prime}=\mathbf{Q}_{E}^{\ast}$, noting that $\mathbf{Q}_{E}^{\ast}\subseteq\mathbf{Q}_{\it MI}^{\ast}$ because Assumption 2.1 implies Assumption 3.2.
A.9 Proof of Corollary 3.3
To begin, note by assumption $\mathbf{Q}_{0}(P,\mathbf{Q}_{\it WE,M}^{\ast})\neq\emptyset$, so there exists $Q\in\mathbf{Q}_{\it WE,M}^{\ast}$ such that $P=QT^{-1}$. We then have $Q\{Z=z\}=P\{Z=z\}$ for all $z\in\mathcal{Z}$, and because $Q$ satisfies Assumption 3.3, we know (2) holds. Let $Q_{1}$ denote the marginal distribution of $((Y_{d}:d\in\mathcal{D}),(D_{z}:z\in\mathcal{Z}))$ under $Q$, and $Q_{Z}$ denote the marginal distribution of $Z$ under $Q$, and define $\tilde{Q}=Q_{1}\times Q_{Z}$. Then, $\tilde{Q}$ satisfies Assumption 2.1 by construction, and it satisfies Assumption 2.2 because the marginal distribution of $(D_{z}:z\in\mathcal{Z})$ under $\tilde{Q}$ is the same as that under $Q$. In summary, $\tilde{Q}\in\mathbf{Q}_{E,M}^{\ast}$. Furthermore, $P=\tilde{Q}T^{-1}$ because (1) $\tilde{Q}\{Z=z\}=Q\{Z=z\}=P\{Z=z\}$ for all $z\in\mathcal{Z}$; and (2) $\tilde{Q}\{Y_{d}=y,D_{z}=d\}=Q\{Y_{d}=y,D_{z}=d\}$ for all $d\in\mathcal{D}$ and $z\in\mathcal{Z}$, and hence (2) is still satisfied. As a result, we know $\mathbf{Q}_{0}(P,\mathbf{Q}_{E,M}^{\ast})\neq\emptyset$.
Next, take $\mathbf{Q}=\mathbf{Q}_{E,M}^{\ast}$ and $\mathbf{Q}^{\prime}=\mathbf{Q}_{\it WE,M}^{\ast}$, and note that $\mathbf{Q}_{E,M}^{\ast}\subseteq\mathbf{Q}_{\it WE,M}^{\ast}\subseteq\mathbf{Q}_{\it MI}^{\ast}$ because Assumption 2.1 implies Assumption 3.3, which in turn implies Assumption 3.2. Because we know $\mathbf{Q}_{0}(P,\mathbf{Q}_{E,M}^{\ast})\neq\emptyset$ from the previous paragraph, we then obtain from Theorem 3.3 that
$\Theta_{0}(P,\mathbf{Q}_{E,M}^{\ast})=\Theta_{0}(P,\mathbf{Q}_{\it WE,M}^{\ast})=\Theta_{0}(P,\mathbf{Q}_{\it MI}^{\ast})~.$
Similarly, taking $\mathbf{Q}^{\prime}=\mathbf{Q}_{\it WE}^{\ast}$, we have
$\Theta_{0}(P,\mathbf{Q}_{E,M}^{\ast})=\Theta_{0}(P,\mathbf{Q}_{\it WE}^{\ast})=\Theta_{0}(P,\mathbf{Q}_{\it MI}^{\ast})~.$
The desired conclusion now follows.
@@ Appendix B Details of Examples
We show that, except in the special case where the treatment and the instrument are both binary, the strict one-to-one targeting assumption of lee2023treatment with one or more targeted treatments implies that (12) does not hold for some treatments. To see this, suppose that the strict one-to-one targeting assumption of lee2023treatment holds with $|\mathcal{D}^{\dagger}|\geq 1$. For each $d\in\mathcal{D}^{\dagger}$, their assumptions include that there exists some $z^{\dagger}(d)$ and some $\overline{U}(d)$, $\underline{U}(d)$ with $\overline{U}(d)>\underline{U}(d)$ such that
@@
g(z,d)>g(z(d),d) for all $d$ and $z,d^{\prime})>g(z^{\ast}($$d^{\prime})~~\text${ for all $^{\prime}\neq d\text${ and $\neq z^{\ast}$(d)~. (21)
@@
We have shown (12) does not hold for some treatments when either $Z$ or $D$ takes at least three values. Now suppose $|\mathcal{D}|=|\mathcal{Z}|=2$. Let $D=0$ denote the nontargeted treatment and $D=1$ the targeted treatment, and let $z^{\dagger}(1)=1$. Consider $z^{\ast}(0)=0$ and $z^{\ast}(1)=1$. Then evaluating (12) at either $d=0$ or $d=1$ results in $\overline{U}(1)>\underline{U}(1),$ and thus (12) holds when $|\mathcal{D}|=|\mathcal{Z}|=2$. We conclude that the strict one-to-one targeting of lee2023treatment implies that (12) does not hold for some $d\in\mathcal{D}$ except in the special case where $|\mathcal{D}|=|\mathcal{Z}|=2$.
We now show that the strict one-to-one targeting of lee2023treatment implies that Assumption 2.2 holds when $|\mathcal{Z}|>|\mathcal{D}^{\dagger}|$. Let $\mathcal{Z}^{\dagger}\subseteq\mathcal{Z}$ denote the set of instruments that target some treatment,
@@
Their strict one-to-one targeting assumption combined with $|\mathcal{Z}|>|\mathcal{D}^{\dagger}|$ implies that there are values of the instrument that do not target any treatment; in other words, $\mathcal{Z}^{\dagger}\subsetneqq\mathcal{Z}$. Following lee2023treatment, we label the treatment that is known not to be targeted as treatment 0, so that $g(z,0)=\underline{U}(0)$ for all $z\in\mathcal{Z}$, and impose their normalization that $\underline{U}(0)=0$. Consider (4) for $d=0$. Note that
@@
Thus, the event $\{D_{z^{\ast}(0)}=d^{\ast},~D_{z^{\prime}}=0\}$ either requires (22) and (23) to jointly hold, which is a contradiction since $\overline{U}(d^{\ast})>\underline{U}(d^{\ast})$ , or requires (22) and (24) to jointly hold, which is a zero probability event given our assumption that the distribution of $(U_{d}:d\in\mathcal{D})$ is absolutely continuous w.r.t. Lebesgue measure. Thus $Q\{D_{z^{\ast}(0)}\neq 0,~D_{z^{\prime}}=0~\mbox{for some}~z^{\prime}\neq z^{\ast}(0)\}$ is a probability of a finite union of zero probability events, and thus, by Boole’s inequality, equals zero so that (4) holds for $d=0$. A parallel argument shows that (4) holds for any non-targeted treatment, and related argument shows that (4) holds for any targeted treatment. Thus, under the strict one-to-one targeting of lee2023treatment, when there are more values of the instrument than targeted treatments, Assumption 2.2 holds even though (12) is violated for some $d\in\mathcal{D}$.
@@
In contrast, the identified set that follows from imposing Assumption 2.1 alone, $\Theta_{0}(P,\mathbf{Q}_{E}^{\ast})$, is shown by balke1997bounds to be
@@
Suppose $\mathcal{Y}=\{0,1\}$, $\mathcal{D}=\{0,1,2\}$, and $\mathcal{Z}=\{0,1,2\}$. Then, the linear program approach in balke1993nonparametric; balke1997bounds leads to the following identified set for $\mathbb{E}_{Q}[Y_{1}]=Q\{Y_{1}=1\}$ relative to $\mathbf{Q}$ being defined as in Example 5.2:
@@
Consider the ARUM defined by (11) with strict one-to-one targeting and $|\mathcal{D}|=3$, $|\mathcal{Z}|=|\mathcal{D}^{\dagger}|=2$. In this case, there are two targeted treatments and one non-targeted treatment. Following lee2023treatment, label that non-targeted treatment as treatment 0 and impose the normalization that $g(z,0)=0$ for all $z\in\mathcal{Z}$. Label $Z=0$ as the instrument value that targets treatment $1$ and label $Z=1$ as the instrument value that targets treatment $2$, so that (20) holds for $d=1,2$ for some $\overline{U}(d),\underline{U}(d)$ with $\overline{U}(d)>\underline{U}(d)$ and with $z^{\dagger}(1)=0$, $z^{\dagger}(2)=1$. Let $\mathbf{Q}$ denote the set of all distributions for which $(D_{z}:z\in\mathcal{Z})$ is determined by (11) with these restrictions and additionally imposing that the support of $(U_{0},U_{1},U_{2})$ is $\Re^{3}$. Recall $(U_{0},U_{1},U_{2})\perp\!\!\!\perp Z$ by assumption.
@@
and a similar characterization holds for $D_{z}=1$. See Figure 1, which is taken from Figure 1 in lee2023treatment.
@@
We now show that there does not exist a value of $z^{\ast}(0)$ such that (4) holds for the non-targeted treatment, treatment 0. Suppose $z^{\ast}(0)=0$. Then
@@
where the last line is using that the support of the distribution of $(U_{10},U_{20})=\Re^{2}$ by assumption and that strict targeting of treatment $1$ requires $\underline{U}(1)>-\overline{U}(1)$. Thus (4) cannot hold for $d=0$ with $z^{\ast}(0)=0$. A parallel argument shows that (4) cannot hold for $d=0$ with $z^{\ast}(0)=1$.
@@
We now derive the identified set for the average potential outcome of the non-targeted treatment. First, note that $g(0,1)=-\overline{U}(1)<-\underline{U}(1)=-g(1,1)$ and $g(0,2)=-\underline{U}(2)>-\overline{U}(2)=-g(1,2)$. Therefore, it can be verified from Figure 2 that for all $Q\in\mathbf{Q}$,
@@
Let $\mathbf{Q}^{\prime}$ denote the set of all distributions that satisfies (30). Note that all $Q\in\mathbf{Q}$ satisfies (30), so $\mathbf{Q}\subseteq\mathbf{Q}^{\prime}$. On the other hand, by assigning appropriate probabilities to each set in the partition in Figure 2, we immediately see that each $Q\in\mathbf{Q}^{\prime}$ can be rationalized by a $Q\in\mathbf{Q}$. Therefore, $\mathbf{Q}=\mathbf{Q}^{\prime}$. Using linear programming as in balke1993nonparametric; balke1997bounds, we obtain the following identified set for $\mathbb{E}_{Q}[Y_{0}]=Q\{Y_{0}=1\}$ relative to $\mathbf{Q}$:
@@
Finally, we show that there exists a $P$ for which $\mathbf{Q}_{0}(P,\mathbf{Q})\neq\emptyset$ and (31) is strictly smaller than (6), so that $\Theta_{0}(P,\mathbf{Q})$ is not given by (6). We do so by providing a numerical example. Consider the $P$ specified in Table 8 and the $Q_{\text{arum},\text{min}}$ and $Q_{\text{arum},\text{max}}$ specified in Tables 9 and 10 respectively, where we write $q(y_{0}y_{1}y_{2},d_{0}d_{1})=Q\{Y_{d}=y_{d},D_{z}=d_{z},~(d,z)\in\mathcal{D}\times\mathcal{Z}\}$ and omit any $q(\cdot)=0$. One can check that both $Q_{\text{arum},\text{min}}$ and $Q_{\text{arum},\text{max}}$ rationalize $P$ and satisfy Assumption 2.1. One can further check that both $Q_{\text{arum},\text{min}}$ and $Q_{\text{arum},\text{max}}$ satisfy the restriction in (30), so that $\mathbf{Q}_{0}(P,\mathbf{Q})\neq\emptyset$. Evaluating (31) at $P$ gives the identified set for $\mathbb{E}_{Q}[Y_{0}]$ relative to $\mathbf{Q}$ as $[0.2518,0.8167]$. One can further check that the two endpoints are attained by $\mathbb{E}_{Q_{\text{arum},\text{min}}}[Y_{0}]=0.2518$ and $\mathbb{E}_{Q_{\text{arum},\text{max}}}[Y_{0}]=0.8167$. On the other hand, if one evaluates (6) by setting $z^{\ast}(0)=0,1$ at the same $P$, the resulting bounds for $\mathbb{E}_{Q}[Y_{0}]$ equal $p_{10|0},1-p_{00|0}]=[0.2518,0.8937]$ and $p_{10|1},1-p_{00|1}]=[0.2372,0.8167]$ respectively. In both cases, (31) is strictly contained in (6). Hence, $\Theta_{0}(P,\mathbf{Q})$ is not given by (6).
@@
kline2016evaluating considers an RCT with a “close substitute” to study the effects of preschooling on educational outcomes. In their setting, $D\in\mathcal{D}=\{0,1,2\}$, where $D=0$ denotes home care (no preschool), $D=2$ denotes a preschool program called Head Start, and $D=1$ denotes preschools other than Head Start, namely the close substitute. Let $Z\in\mathcal{Z}=\{0,1\}$ denote an indicator variable for an offer to attend Head Start. Assumption 2.1 holds because $Z$ is randomly assigned. kline2016evaluating impose the restriction that
@@
Note in this example Assumption 2.2 still holds although $|\mathcal{Z}|<|\mathcal{D}|$. See bai2025sharp for results on the sharp testable implications of the assumptions for this example and Example C.3.
@@
kirkeboen2016field study the effects of fields of study on earnings. In their setting, $\mathcal{D}=\{0,1,2\}$ represent three fields of study, ordered by their (soft) admission cutoffs from the lowest to the highest. The instrument is $Z\in\{0,1,2\}$, with $Z=1$ when the student crosses the (soft) admission cutoff for field 1, $Z=2$ when the student crosses the (soft) admission cutoff for field 2, and $Z=0$ otherwise. The authors assume that $Z$ is exogenous in the sense that $Q$ satisfies Assumption 2.1 and impose the following monotonicity conditions: